OpenStudy (anonymous):

solve each system by elimination Please help me...Reward given :) 2x+18y=22 -x-9y=-11

OpenStudy (osanseviero):

(2,1)

OpenStudy (anonymous):

Infinitely Many Solutions

OpenStudy (anonymous):

what does that meanm

OpenStudy (anonymous):

it means its the same line but written differently

OpenStudy (anonymous):

it isn't infinite

hero (hero):

It means both equations have the same solutions

OpenStudy (anonymous):

they are on the same line therefore infinite solutions

OpenStudy (anonymous):

-x-9y=-11 x+9y=11 2x+18y=22 same as the first equation

OpenStudy (anonymous):

two unknowns, two equations. two lines.

OpenStudy (anonymous):

GIRL you need to get some help because you obvi don't understand this

OpenStudy (anonymous):

different coefficients

hero (hero):

A bunch of confusing responses from everyone doesn't help

OpenStudy (anonymous):

but she aint even trying just keep bombarding us with questions and then jumping to the next one

OpenStudy (anonymous):

ok if u have 2 lines that are parallel but not the same then there is no solution. if u have to lines that are the same then u have infintely many solutions and if u have 2 lines that are not parallel and not the same then it has1 unique solution. this is where they intersect

OpenStudy (anonymous):

i dont mind helping with homework as long as im not doing it all, they're learning some too

OpenStudy (anonymous):

They are the same line, so infinite would be the correct answer.